A chemist mixes two solutions. One is 24% acid (the rest is water), and the other is 41% acid. About how much of each solution does she need to produce 50 gallons of a solution that is 31% acid? Answer to the nearest hudredth of a gallon.
Answer by Paul(988)

24x+41(50-x)=31(50)
24x+2050-41x=1550
-17x=-500
x=20.41
50-29.41=20.59

Hence, for 24% there is 29.41G needed and for 41% there is 20.59G needed.
PAUL.